# Decomposable Submodular Maximization in Federated Setting

### v1

@@ 6 Conclusions and Future Work
We present FedCG, the first algorithm for decomposable submodular maximization in federated setting under matroid constraints. FedCG is based on the continuous greedy algorithm and achieves the best possible approximation factor i.e., $1-1/\mathrm{e}$ under mild assumptions even faced with client selection and low communication rounds. Additionally, we introduce a new federated framework for discrete problems. Our work leads to many interesting directions for future work, such as providing stronger privacy guarantee using differentially private methods developed for submodular maximization (Mitrovic et al., 2017; Rafiey and Yoshida, 2020; Chaturvedi et al., 2021), finding a lower bound on the bit complexity, and improving the additive error of our algorithms.
Acknowledgement
@@ Appendix A Preliminary results and proof of convergence for full participation
At the heart of the analyses for the approximation ratio of the Centralized Continuous Greedy is to show that at each iteration the algorithm reduces the gap to the optimal solution by a significant amount [Călinescu et al., 2011]. We follow the same general idea although there are several subtleties that we should address. Mainly, it is required to compare the improvement we obtain by taking direction $\mathbf{v}$ versus the improvement obtained by taking direction $\sum_{i=1}^{N}p_{i}\mathbf{v}_{i}$ where $\mathbf{v}=\operatornamewithlimits{argmax}_{\mathbf{y}\in\mathcal{P}}\langle\mathbf{y},\nabla\widehat{F}(\mathbf{x})\rangle$ and $\mathbf{v}_{i}=\operatornamewithlimits{argmax}_{\mathbf{y}\in\mathcal{P}}\langle\mathbf{y},\nabla\widehat{f}_{i}(\mathbf{x})\rangle$. To establish a comparison between the two, we need several intermediate lemmas.
@@
A.1 Putting everything together, proof of Theorem 3.2.
@@
where $\mathbf{H}_{\widehat{F}}$ is the Hessian matrix i.e., the second derivative matrix. We provide a lower bound on each entry of $\mathbf{H}_{\widehat{F}}$. By the result of [Călinescu et al., 2011] and definition of the multilinear extension:
@@ B.1 Bounding the variance
Bounding the variance is at the core of our proof of convergence. Recall the definition of $\gamma_{t}$ in Assumption 3.1 which plays a pivotal role in bounding the variance.
@@
where in the last inequality we used the fact that for each $\mathbf{v}_{i}^{(t)}$ we have $\mathbf{v}_{i}^{(t)}\in\{0,1\}^{n}$ and $|\mathbf{v}_{i}^{(t)}|_{1}=r$, and Assumption 3.1 that together yield $|\langle\mathbf{v}_{i},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle|\leq r\gamma_{t}$.
@@
By Assumption 3.1 we know that $|\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})|_{\infty}\leq\gamma_{t}$ and $|\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})|_{\infty}\leq\gamma_{t}$. Therefore, $|\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})-\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})|_{\infty}\leq 2\gamma_{t}$. Knowing that both $\mathbf{v}_{i}^{(t)},\mathbf{s}\in\{0,1\}^{n}$ and $|\mathbf{v}_{i}^{(t)}|_{1}=|\mathbf{s}|_{1}=r$ yields
@@
B.2 Putting everything together, proof of Theorem 3.5.
@@ C.1 Bounding the variance
Bounding the variance is at the core of our proof of convergence. Assumption 3.1 plays a pivotal role in bounding the variance. The main difficulty and difference between this proof and the proof of Lemma 3.4 is that firstly the variance is effected by the divergence caused by local steps, and secondly $\widetilde{\Delta}_{i}^{(t+\tau)}$ may not be integral vectors and it could potentially have $O(n)$ nonzero entries.
@@
$\leq\sum_{i=1}^{N}p_{i}\left(\frac{1}{\tau}\sum_{j=0}^{\tau-1}r\gamma_{t}\right)+(\sigma r+L_{t}r^{1.5})$ (Assumption 3.1)
@@
By Assumption 3.1 we know that $|\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})|_{\infty}\leq\gamma_{t}$ and $|\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})|_{\infty}\leq\gamma_{t}$. Therefore, $|\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})-\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})|_{\infty}\leq 2\gamma_{t}$. This is used to upper bound the first term in (90). In order to bound the second term in (90) we appeal to equation (72) in Case 1. (Note that there are two cases two consider here because of the absolute value, however the argument is similar and we argue about one case.) Let $\mathbf{1}$ be the all one vector of length $n$, then
@@
$\left(1-(1-\eta)^{T/\tau}\right){\texttt{OPT}}$ $\leq\widehat{F}(\mathbf{x}^{(T/\tau)})+\eta\left(r\sum_{t\in I_{\tau}}\gamma_{t}+\frac{6r\sum_{t\in I_{\tau}}\gamma_{t}+2(\sigma rT/\tau+\sum_{t\in I_{\tau}}L_{t}r^{1.5})}{\sqrt{K\tau\delta/T}}+2(\sigma rT/\tau+r^{1.5}\sum_{t\in I_{\tau}}L_{t})\right)$ (109)
@@ C.3 Gradient estimation
In terms of computation cost on clients’ devices, we point out that the definition of multilinear extension involves summing over all subsets $S$ of $E$. There are $2^{|E|}$ such subsets, thus even computing $\widehat{g}(\mathbf{x})$ for a single $\mathbf{x}$ could take exponential time. However, we can randomly sample $m$ subsets $R_{1},\dots,R_{m}$ of $E$ according to $\mathbf{x}$. Then a simple application of Chernoff’s bound shows for any multilinear extension $\widehat{g}$ and $\mathbf{x}$, $\left|\frac{1}{m}\sum_{i=1}^{m}g(R_{i})-\widehat{g}(\mathbf{x})\right|\leq\sigma\max_{S}g(S)$ [Vondrák, 2008, Călinescu et al., 2011] with probability at least $1-\mathrm{e}^{-m\sigma^{2}/4}$. Observe that $\frac{\partial\widehat{g}}{\partial\mathbf{x}(e)}=\mathbb{E}[g(R\cup\{e\})]-\mathbb{E}[g(R)]$ where $R$ is a random subset of $E\setminus\{e\}$ sampled according to $\mathbf{x}$. Hence, by a similar argument, with $m$ random samples, we can compute an $\sigma$-approximation of $\nabla\widehat{g}(\mathbf{x})$ with $1-\mathrm{e}^{-m\sigma^{2}/4}$ probability. By suitably choosing $m$ as in Algorithm 2, we can assume the gradients are estimated within $(1\pm\sigma)$ accuracy with high probability.
@@
Appendix D Discrete examples
@@
D.0.1 Facility Location Problem
@@
Here we discuss that each client can compute its corresponding importance factor in a federated setting and using a secure aggregator without sharing its data with other clients. Because of the simple nature of importance factors for this problem, one can easily verify that in Algorithm 4 every client correctly computes its own importance factor. Furthermore, Algorithm 4 requires only two communication rounds.
@@
D.0.2 Maximum Coverage Problem
@@
Here we discuss that each client can compute its corresponding importance factor in a federated setting and using a secure aggregator without sharing which group she belongs to. Because of the simple nature of importance factors for this problem, one can easily verify that in Algorithm 5 every client correctly computes its own importance factor. Furthermore, Algorithm 5 requires only two communication rounds.
Theorem D.3.
Remark D.4.
We point out that a simple algorithm where at each round clients share their membership with the server using SecAgg protocols is applicable here. However, such algorithm requires full client participation at each round. This can be resolved by sampling sufficiently many clients at each round. However, the number of sampled clients as well as the approximation factor would depend on $N$.
Having $w_{i}$ on hand we can execute Algorithm 3 for Max Coverage problem. Note that, in this problem we are dealing with a uniform matroid of rank $k$.
Theorem D.5.
$(1-1/\mathrm{e}-\varepsilon)\texttt{OPT}\leq F(S)$
Proof.
$\sum_{i=1}^{n}\kappa_{i}\leq\kappa\sum_{i=1}^{n}w_{i}\leq\tilde{O}(kn/\varepsilon^{2})\sum_{i=1}^{N}w_{i}=\tilde{O}(kn/\varepsilon^{2})\sum_{i=1}^{N}\max\limits_{\begin{subarray}{c}G_{a}\in E,C_{i}\in G_{a}\end{subarray}}\frac{1}{|G_{a}|}\leq\tilde{O}(kn/\varepsilon^{2})\sum_{j=1}^{|E|}\frac{|G_{j}|}{|G_{j}|}=\tilde{O}(kn^{2}/\varepsilon^{2})$

### v2

@@ 6 Conclusions and Future Work
We present FedCG, the first algorithm for decomposable submodular maximization in federated setting under matroid constraints. FedCG is based on the continuous greedy algorithm and achieves the best possible approximation factor i.e., $1-1/\mathrm{e}$ under mild assumptions even faced with client selection and low communication rounds. Additionally, we introduce a new federated framework for discrete problems. Our work leads to many interesting directions for future work, such as providing stronger privacy guarantee using differentially private methods (Mitrovic et al., 2017; Rafiey & Yoshida, 2020; Chaturvedi et al., 2021), finding a lower bound on the bit complexity, and improving the additive errors of our algorithms.
Acknowledgements
@@
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Preliminary Results and Proof of Convergence for Full Participation
At the heart of the analyses for the approximation ratio of the Centralized Continuous Greedy is to show that at each iteration the algorithm reduces the gap to the optimal solution by a significant amount (Călinescu et al., 2011). We follow the same general idea although there are several subtleties that we should address. Mainly, it is required to compare the improvement we obtain by taking direction $\mathbf{v}$ versus the improvement obtained by taking direction $\sum_{i=1}^{N}p_{i}\mathbf{v}_{i}$ where $\mathbf{v}=\operatornamewithlimits{argmax}_{\mathbf{y}\in\mathcal{P}}\langle\mathbf{y},\nabla\widehat{F}(\mathbf{x})\rangle$ and $\mathbf{v}_{i}=\operatornamewithlimits{argmax}_{\mathbf{y}\in\mathcal{P}}\langle\mathbf{y},\nabla\widehat{f}_{i}(\mathbf{x})\rangle$. To establish a comparison between the two, we need several intermediate lemmas.
@@
A.1 Putting Everything Together: Proof of Theorem 3.2
@@
where $\mathbf{H}_{\widehat{F}}$ is the Hessian matrix i.e., the second derivative matrix. We provide a lower bound on each entry of $\mathbf{H}_{\widehat{F}}$. By the result of (Călinescu et al., 2011) and definition of the multilinear extension:
@@
A.2 Upper Bound for $\gamma_{t}$ in 3.1
Inequality (5) in 3.1 can be derived using monotonicity and submodularity as follows. For simplicity we drop $t$.
$|\nabla\widehat{f}_{i}(\mathbf{x})-\nabla\widehat{F}(\mathbf{x})|_{\infty}$ $=\max_{e\in E}|[\nabla\widehat{f}_{i}(\mathbf{x})]_{e}-[\nabla\widehat{F}(\mathbf{x})]_{e}|$
$\leq\max_{e\in E}[\nabla\widehat{f}_{i}(\mathbf{x})]_{e}+[\nabla\widehat{F}(\mathbf{x})]_{e}$ (by monotonicity and submodularity: $[\nabla\widehat{F}(\mathbf{x})]_{e},[\nabla\widehat{f}_{i}(\mathbf{x})]_{e}\geq 0$)
$\leq\max_{e\in E}\max_{i}[\nabla\widehat{f}_{i}(\mathbf{x})]_{e}+[\nabla\widehat{f}_{i}(\mathbf{x})]_{e}$ (using $F=\sum_{i}p_{i}f_{i}$ and $\sum_{i}p_{i}=1$)
$=2\max_{e\in E}\max_{i}[\nabla\widehat{f}_{i}(\mathbf{x})]_{e}$
$=2\max_{i}\max_{e\in E}[\nabla\widehat{f}_{i}(\mathbf{x})]_{e}$
$=2\max_{i}m_{f_{i}}$
@@ B.1 Bounding the Variance
Bounding the variance is at the core of our proof of convergence. Recall the definition of $\gamma_{t}$ in 3.1 which plays a pivotal role in bounding the variance.
@@
where in the last inequality we used the fact that for each $\mathbf{v}_{i}^{(t)}$ we have $\mathbf{v}_{i}^{(t)}\in\{0,1\}^{n}$ and $|\mathbf{v}_{i}^{(t)}|_{1}=r$, and 3.1 together yield $|\langle\mathbf{v}_{i},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle|\leq r\gamma_{t}$.
@@
By 3.1 we know that $|\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})|_{\infty}\leq\gamma_{t}$ and $|\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})|_{\infty}\leq\gamma_{t}$. Therefore, $|\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})-\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})|_{\infty}\leq 2\gamma_{t}$. Knowing that both $\mathbf{v}_{i}^{(t)},\mathbf{s}\in\{0,1\}^{n}$ and $|\mathbf{v}_{i}^{(t)}|_{1}=|\mathbf{s}|_{1}=r$ yields
@@
B.2 Putting Everything Together: Proof of Theorem 3.5
@@ C.1 Bounding the variance
Bounding the variance is at the core of our proof of convergence. 3.1 plays a pivotal role in bounding the variance. The main difficulty and difference between this proof and the proof of Lemma 3.4 is that firstly the variance is effected by the divergence caused by local steps, and secondly $\widetilde{\Delta}_{i}^{(t+\tau)}$ may not be integral vectors and it could potentially have $O(n)$ nonzero entries.
@@
$\leq\sum_{i=1}^{N}p_{i}\left(\frac{1}{\tau}\sum_{j=0}^{\tau-1}r\gamma_{t}\right)+(\sigma r+L_{t}r^{1.5})$ (3.1)
@@
By 3.1 we know that $|\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})|_{\infty}\leq\gamma_{t}$ and $|\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})|_{\infty}\leq\gamma_{t}$. Therefore, $|\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})-\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})|_{\infty}\leq 2\gamma_{t}$. This is used to upper bound the first term in (87). In order to bound the second term in (87) we appeal to equation (69) in Case 1. (Note that there are two cases two consider here because of the absolute value, however the argument is similar and we argue about one case.) Let $\mathbf{1}$ be the all one vector of length $n$, then
@@
$\left(1-(1-\eta)^{T/\tau}\right){\texttt{OPT}}$ $\leq\widehat{F}(\mathbf{x}^{(T/\tau)})$ (106)
$+\eta\left(r\sum_{t\in I_{\tau}}\gamma_{t}+\frac{6r\sum_{t\in I_{\tau}}\gamma_{t}+2(\sigma rT/\tau+\sum_{t\in I_{\tau}}L_{t}r^{1.5})}{\sqrt{K\tau\delta/T}}+2(\sigma rT/\tau+r^{1.5}\sum_{t\in I_{\tau}}L_{t})\right)$ (107)
@@ C.3 Gradient Estimation
In terms of computation cost on clients’ devices, we point out that the definition of multilinear extension involves summing over all subsets $S$ of $E$. There are $2^{|E|}$ such subsets, thus even computing $\widehat{g}(\mathbf{x})$ for a single $\mathbf{x}$ could take exponential time. However, we can randomly sample $m$ subsets $R_{1},\dots,R_{m}$ of $E$ according to $\mathbf{x}$. Then a simple application of Chernoff’s bound shows for any multilinear extension $\widehat{g}$ and $\mathbf{x}$, $\left|\frac{1}{m}\sum_{i=1}^{m}g(R_{i})-\widehat{g}(\mathbf{x})\right|\leq\sigma\max_{S}g(S)$ (Vondrák, 2008; Călinescu et al., 2011) with probability at least $1-\mathrm{e}^{-m\sigma^{2}/4}$. Observe that $\frac{\partial\widehat{g}}{\partial\mathbf{x}(e)}=\mathbb{E}[g(R\cup\{e\})]-\mathbb{E}[g(R)]$ where $R$ is a random subset of $E\setminus\{e\}$ sampled according to $\mathbf{x}$. Hence, by a similar argument, with $m$ random samples, we can compute an $\sigma$-approximation of $\nabla\widehat{g}(\mathbf{x})$ with $1-\mathrm{e}^{-m\sigma^{2}/4}$ probability. By suitably choosing $m$ as in Algorithm 2, we can assume the gradients are estimated within $(1\pm\sigma)$ accuracy with high probability.
@@
Appendix D Discrete Algorithm in Federated Setting: Examples
@@
D.1 Facility Location Problem
@@
In many applications computing importance factors is straightforward. Let us elaborate on this with an real-world application.
Movie recommendation system.
Consider a movie recommendation application (Stan et al., 2017) where each client $i$ has a user-specific utility function $f_{i}$ to evaluate sets of movies. The global task is to find a set of $k$ movies that are most satisfactory to all the clients. An example is the MovieLens dataset consisting of 1 million ratings by $N=6041$ clients for $|E|=n=4000$ movies. It is in the interest of clients that we respect their privacy and they are reluctant to share their rating with a central server and other clients. We consider a well motivated objective function. Let $r_{i,j}$ denote the rating of client $i$ for movie $j$ (if such a rating does not exist set $r_{i,j}=0$). We associate to each client $i$ a facility location objective function $f_{i}(S)=\max_{j\in S}r(i,j)$ where $S\subseteq E$ is a subset of movies. The servers objective is $\max_{S\subseteq E,|S|\leq k}\frac{1}{N}\sum f_{i}(S)$. In this example, the average rating of each movie is publicly available. That is $F(\{j\})=1/N\sum_{i=1}^{N}f_{i}(\{j\})$ for each movie $j$ is publicly available. Hence, it is straightforward for each client to compute its own importance factor.
@@
It is straightforward to see each client can compute its corresponding importance factor in a federated setting and using a secure aggregator without sharing its data with other clients. Each client $i$ sends a vector $(c(i,1),\dots,c(i,n))$ to the server and by simply summing up these vectors the server has a histogram over facilities. This histogram is then broadcasts to the clients where they can compute their own importance factor; see Algorithm 4). Furthermore, Algorithm 4 requires only two communication rounds.
@@
D.2 Maximum Coverage Problem
@@
Having a histogram over the group sizes suffices for computing the importance factors. Similar to Facility Location the importance factors can be computed in federated setting by having the membership histograms over the groups; see Algorithm 5.
@@
Theorem D.3.
Remark D.4.
We point out that a simple algorithm where at each round clients share their membership with the server using SecAgg protocols is applicable here. However, such algorithm requires full client participation at each round. This can be resolved by sampling sufficiently many clients at each round.
Having $w_{i}$, we can execute Algorithm 3 for Max Coverage problem. Note that, in this problem we are dealing with a uniform matroid of rank $k$.
Theorem D.5.
$(1-1/\mathrm{e}-\varepsilon)\texttt{OPT}\leq F(S)$
Proof.
$\sum_{i=1}^{n}\kappa_{i}\leq\kappa\sum_{i=1}^{n}w_{i}\leq\tilde{O}(kn/\varepsilon^{2})\sum_{i=1}^{N}w_{i}=\tilde{O}(kn/\varepsilon^{2})\sum_{i=1}^{N}\max\limits_{\begin{subarray}{c}G_{a}\in E,C_{i}\in G_{a}\end{subarray}}\frac{1}{|G_{a}|}\leq\tilde{O}(kn/\varepsilon^{2})\sum_{j=1}^{|E|}\frac{|G_{j}|}{|G_{j}|}=\tilde{O}(kn^{2}/\varepsilon^{2})\qed$